OpenStudy (anonymous):

3(3x - 2y) = -10 (2 + 5y) = 6, solve for x?

OpenStudy (radar):

I am thinking you will need to solve for y first. Using the 2nd equation solve for y. 2 + 5y = 6 5y = 6 - 2 = 4 5y = 4 y= 4/5 can you take it from there?

OpenStudy (anonymous):

Sadly no, I wish I understand the problems fully

OpenStudy (radar):

did you follow how y was solved from the 2nd equation?

OpenStudy (anonymous):

Sort of yes actually! Moving onto 3(2x - 2) = -10, we would multiply 3 and 3x together, then subtract 2y correct?

OpenStudy (radar):

That was the first equation, You must solve for y first. That is what I did in my first post. y was solved using the 2nd equation. Only simple algebra is needed. The first line is just rewriting the equation as it was written in your problem. The second line 2 was subtracted from both sides. 3rd line is simply writing the results. 4th line is dividing both sides by 5. Do you follow how I solved for y??

OpenStudy (radar):

Think about it and substitute the y value into the y in the first equation and solve for x which is the mission for the problem. I gotta run now, good luck with the problem.